# MSPCaps: A Multi-Scale Patchify Capsule Network with Cross-Agreement Routing for Visual Recognition

### v1

@@ 32 $\times$ 32? 16 $\times$ 16? 8 $\times$ 8? Params CIFAR-10
To fully explore the potential of our architecture under parameter-sufficient settings, we compare the full-sized model–MSPCaps-L with other prominent and deeper baselines. As shown in Tab. 3, MSPCaps-L exhibits the best performance on all datasets, specifically achieving accuracies of 99.73% on MNIST, 95.05% on FashionMNIST, and a notable 92.88% on CIFAR-10. In the category of deep CapsNet, our MSPCaps-L demonstrates a significant performance gain compared to DeepCaps by 1.87% on CIFAR-10. This can be attributed to the powerful multi-scale architecture. The improvement is further enhanced when compared to the smaller models, i.e., CapsNet and AR-CapsNet. The increase in parameter counts for our MSPCaps-L results in a significant 3.48% and 2.77% accuracy improvement on CIFAR-10, respectively. Furthermore, compared to the multi-scale based RS-CapsNet, our model outperforms with a significant gap of 3.07% on CIFAR-10 which demonstrates the superiority of our architecture. These results validate our model’s scalability and the ability to convert increased parameter budget into superior feature representation.
@@
Contribution Analysis of Multi-Scale Features. To demonstrate the contribution of multi-scale capsules for final prediction, we compare our full MSPCaps-T model against several tiny ablated variants on CIFAR-10, as shown in Tab 4. These variants use capsules generated from specific feature map scales ($32\times 32$, $16\times 16$, and $8\times 8$). In the single-scale experiments, we input two identical sets of capsules, adopting the same capsule dimension configuration as the first CAR block in MSPCaps-T. The results affirm the effectiveness of our multi-scale design, with the full model achieving a top accuracy of 88.71%. The $8\times 8$ feature map acts as the primary contributor, establishing the most robust semantic foundation with its global contextual view compared to other ablated variants. However, this global view lacks information on fine details, which is addressed by the high-resolution $32\times 32$ map. By supplying essential fine-grained details for voting, it helps the model resolve ambiguities in classification. Although this detail-oriented contribution is insufficient by itself, it is still indispensable for refining the model’s predictions. Overall, our full MSPCaps-T model excels in its ability to synthesize hierarchical contributions by marrying the crucial details from the fine scale with the foundational understanding from the coarse scale.
@@
Table 6: Ablation study of weight sharing within CAR.
@@
Quantitative Comparison between DR and CAR. To validate the proposed CAR block, we compare it with the original dynamic routing (DR) [23] on both MSPCaps-T and MSPCaps-L. The ablated variants replace the CAR blocks with DR blocks, where three routing blocks first process the multi-scale features independently, after which their outputs are concatenated and passed to a final, fourth routing block. As shown in Tab. 5, the proposed CAR block demonstrates superior performance and scalability across both model variants. For MSPCaps-T, our CAR variant significantly surpasses DR while reducing parameters by 28.7%. This advantage is further amplified on MSPCaps-L, where our CAR variant achieves the highest accuracy with a 25% parameter reduction compared to DR. These results highlight CAR’s effectiveness in fusing cross-scale capsules for voting, offering consistent gains in both accuracy and parameter efficiency.
Share Weight Mechanism Analysis. Finally, we conduct an ablation study on the effect of weight sharing within the CAR block. The results are shown in Tab. 6. We can observe an interesting trade-off depending on model capacity. Specifically, for the lightweight MSPCaps-T, weight sharing proves to be an effective regularization strategy, enhancing mutual information utilization in parameter-constrained scenarios. In contrast, for the larger MSPCaps-L, disabling weight sharing improves the “part-to-whole" voting process by leveraging its ample parameters.
Table 7: Analysis of different patch sizes for PatchifyCaps.
@@
Patch Size Analysis
To determine the optimal patch size $p$, we evaluate classification accuracy on the CIFAR-10 dataset with patch sizes ranging from $2$ to $4$. As shown in Tab. 7, the lightweight MSPCaps-T model achieves its best performance at $p=4$, while the larger MSPCaps-L model performs optimally at $p=3$. For MSPCaps-T, a patch size of $p=3$ causes uneven division of the feature map, resulting in border feature loss and degraded accuracy. In contrast, MSPCaps-L achieves an optimal parameter balance with $p=3$, leveraging its larger capacity. However, a smaller patch size of $p=2$ introduces excessive parameterization, leading to overfitting and poor generalization. Therefore, to ensure a consistent and efficient architecture, we select $p=4$ as the global setting. This trade-off significantly improves MSPCaps-T’s performance and efficiency, with only a minor accuracy reduction for MSPCaps-L.
Adversarial Robustness Analysis
To demonstrate the robustness of the proposed MSPCaps model, we conduct FGSM [8] and BIM attacks [16] on the CIFAR-10 dataset. For the iterative attack, BIM is performed for 10 steps. The results are shown in Fig. 4. We can observe a clear distinction between our model variants. For MSPCaps-L, it exhibits exceptional stability, maintaining a high level of accuracy even as the attack strength increases. In contrast, constrained by the parameter setting, MSPCaps-T suffers a performance degradation for $\epsilon>0.1$, at which point its classification accuracy falls below that of the original CapsNet.
@@ Conclusion
In this paper, we propose a novel multi-scale patchify capsule network with cross-agreement routing, named MSPCaps. We first introduce MSRB and PatchifyCaps to obtain primary capsules from multi-scale features. Then, to fuse cross-scale capsules effectively for voting, we design a CAR block to adaptively select the most coherent capsules for voting. Experimental results on four datasets demonstrate that MSPCaps outperforms state-of-the-art methods in terms of classification performance and model robustness.
Appendix A Appendix
1. Dataset Statistics
Tab. 8 summarizes the statistics of the four datasets used in our experiments, including the number of data samples, the number of classes, the image resolution, and the preprocessing operations. Specifically, MNIST consists of 70,000 grayscale images of handwritten digits, with 60,000 training samples and 10,000 test samples, each sized $28\times 28$ pixels across 10 classes. FashionMNIST contains 70,000 grayscale images of fashion products, also split into 60,000 training samples and 10,000 test samples spanning 10 clothing categories. Each image in FashionMNIST is sized $28\times 28$ pixels. For SVHN, we use the small version training set, which contains 73,257 training images and 26,032 test images, all in RGB format with a resolution of $32\times 32$ pixels, depicting house numbers from Google Street View across 10 digit classes. CIFAR-10 consists of 60,000 RGB natural images with a resolution of $32\times 32$ pixels, partitioned into 50,000 training samples and 10,000 test samples, evenly distributed across 10 object categories, including animals and vehicles.
For random rotation, we rotate images within a range of $(-15^{\circ},15^{\circ})$ using uniform sampling. For random cropping, we first apply padding to the left, top, right, and bottom edges of each image by 0, 2, 3, and 4 pixels, respectively. Then, a random crop of the original resolution is extracted from the padded image. For normalization, both the training and test sets for MNIST, FashionMNIST, and SVHN are normalized with a mean of 0.5 and a standard deviation of 0.5. For CIFAR-10, we use a channel-wise mean of (0.4914,0.4822,0.4465) and a standard deviation of (0.2023,0.1994,0.2010). More details of the dataset statistics are provided in the source code.
2. Accuracy and Loss Analysis
To analyze the training and testing dynamics, we present the accuracy and loss curves for each model across the datasets. As shown in Fig. 5, the large model (MSPCaps-L) exhibits a faster convergence rate on MNIST compared to the tiny model (MSPCaps-T). However, their final performance is nearly identical, with both converging to approximately 99.7% accuracy. In contrast, Fig. 6 reveals a different trend: the large model not only maintains its superior convergence speed but also achieves a substantial accuracy improvement. It surpasses 90% accuracy within the first 50 epochs, while the tiny model struggles to reach 88% even after 200 epochs. Similar trends can also be observed in Fig. 7 and Fig. 8. A notable shared characteristic is the stability of the training process. Regardless of model size, our method ensures a steady training trajectory with minimal loss fluctuations, demonstrating a significant improvement in stability compared to the conventional dynamic routing [23]. This highlights the robustness of our model’s architecture and training dynamics.
3. Implementation Details
Our implementation incorporates several key configurations to ensure training stability and optimize performance. For weight initialization, the patchify convolution layer adopts Xavier normal distribution [7], while other convolutional layers and projection weights leverage Kaiming initialization [10]. In addition, we employ distinct training settings for different datasets. For the MNIST dataset, the model is trained for 100 epochs, while for the FashionMNIST dataset, we adjust the learning rate to $1\times 10^{-4}$. Regarding regularization, a dropout [24] rate of $0.1$ is applied within the CAR block after computing coupling coefficients. However, this dropout is omitted when training on the SVHN dataset.
Table 8: Statistics of the four datasets.
Dataset Train Samples Test Samples Classes Resolution Preprocessing
MNIST 60,000 10,000 10 $28\times 28$
Resize to $32\times 32$;
Random rotation $\pm 15^{\circ}$;
Normalize
FashionMNIST 60,000 10,000 10 $28\times 28$
Resize to $32\times 32$;
Random horizontal flip;
Random crop with padding;
Normalize
SVHN 73,257 26,032 10 $32\times 32$
Random rotation $\pm 15^{\circ}$;
Random crop with padding;
Normalize
CIFAR-10 50,000 10,000 10 $32\times 32$
Random horizontal flip;
Random crop with padding;
Normalize
(a) MSPCaps-T Accuracy
(b) MSPCaps-T Loss
(c) MSPCaps-L Accuracy
(d) MSPCaps-L Loss
Figure 5: Accuracy and loss curves on the MNIST dataset.
(a) MSPCaps-T Accuracy
(b) MSPCaps-T Loss
(c) MSPCaps-L Accuracy
(d) MSPCaps-L Loss
Figure 6: Accuracy and loss curves on the CIFAR-10 dataset.
(a) MSPCaps-T Accuracy
(b) MSPCaps-T Loss
Figure 7: Accuracy and loss curves for SVHN.
(a) MSPCaps-L Accuracy
(b) MSPCaps-L Loss
Figure 8: Accuracy and loss curves for FashionMNIST.

### v2

@@ 32 $\times$ 32? 16 $\times$ 16? 8 $\times$ 8? Params CIFAR-10
To fully explore the potential of our architecture under parameter-sufficient settings, we compare the full-sized model–MSPCaps-L with other deeper baselines. As shown in Tab. 3, MSPCaps-L exhibits the best performance on all datasets, specifically achieving accuracies of 99.73% on MNIST, 95.05% on FashionMNIST, and a notable 92.88% on CIFAR-10. In the category of deep CapsNet, our MSPCaps-L demonstrates a significant performance gain compared to DeepCaps by 1.87% on CIFAR-10. This can be attributed to the powerful multi-scale architecture. The improvement is further enhanced when compared to the smaller models, i.e., CapsNet and AR-CapsNet. The increase in parameter counts for our MSPCaps-L results in a significant 3.48% and 2.77% accuracy improvement on CIFAR-10, respectively. Furthermore, compared to the multi-scale based RS-CapsNet, our model outperforms with a significant gap of 3.07% on CIFAR-10 which demonstrates the superiority of our architecture. These results validate our model’s scalability and the ability to convert an increased parameter budget into robust feature representation.
@@
Contribution Analysis of Multi-Scale Features. To demonstrate the contribution of multi-scale capsules for final prediction, we compare our full MSPCaps-T model against several tiny ablated variants on CIFAR-10, as shown in Tab 4. These variants use capsules generated from specific feature map scales ($32\times 32$, $16\times 16$, and $8\times 8$). In the single-scale experiments, we input two identical sets of capsules, adopting the same capsule dimension configuration as the first CAR block in MSPCaps-T. The results affirm the effectiveness of our multi-scale design, with the full model achieving a top accuracy of 88.71%. The $8\times 8$ feature map acts as the primary contributor, establishing the most robust semantic foundation with its global contextual view compared to other ablated variants. However, this global view lacks information on fine details, which is addressed by the high-resolution $32\times 32$ map. By supplying essential fine-grained details for voting, it helps the model resolve ambiguities in classification. Although this detail-oriented contribution is insufficient by itself, it is still indispensable for refining the model’s predictions.
@@
Table 6: Ablation study of weight sharing within CAR.
@@
Table 7: Analysis of different patch sizes for PatchifyCaps.
Quantitative Comparison between DR and CAR. To validate the proposed CAR block, we compare it with the original dynamic routing (DR) (Sabour et al. 2017) on both MSPCaps-T and MSPCaps-L. The ablated variants replace the CAR blocks with DR blocks, where three routing blocks first process the multi-scale features independently, after which their outputs are concatenated and passed to a final, fourth routing block. As shown in Tab. 5, the proposed CAR block demonstrates superior performance and scalability across both model variants. For MSPCaps-T, our CAR variant significantly surpasses DR while reducing parameters by 28.7%. This advantage is further amplified on MSPCaps-L, where our CAR variant achieves the highest accuracy with a 25% parameter reduction compared to DR. These results highlight CAR’s effectiveness in fusing cross-scale capsules for voting, offering consistent gains in both accuracy and parameter efficiency.
Share Weight Mechanism Analysis. Finally, we conduct an ablation study on the effect of weight sharing within the CAR block. The results are shown in Tab. 6. We can observe an interesting trade-off depending on model capacity. Specifically, for the lightweight MSPCaps-T, weight sharing proves to be an effective regularization strategy, enhancing mutual information utilization in parameter-constrained scenarios. In contrast, for the larger MSPCaps-L, disabling weight sharing improves the “part-to-whole” voting process by leveraging its ample parameters.
Patch Size Analysis
To determine the optimal patch size $p$, we evaluate classification accuracy on the CIFAR-10 dataset with patch sizes ranging from $2$ to $4$. As shown in Tab. 7, MSPCaps-T achieves its best performance at $p=4$, while the larger MSPCaps-L performs optimally at $p=3$. For MSPCaps-T, a patch size of $p=3$ causes uneven division of the feature map, resulting in border feature loss and degraded accuracy. In contrast, MSPCaps-L achieves an optimal parameter balance with $p=3$, leveraging its larger capacity. However, a smaller patch size of $p=2$ introduces excessive parameterization, leading to overfitting and poor generalization. Therefore, to ensure a consistent and efficient architecture, we select $p=4$ as the global setting. This trade-off significantly improves MSPCaps-T’s performance and efficiency, with only a minor accuracy reduction for MSPCaps-L.
Adversarial Robustness Analysis
To demonstrate the robustness of MSPCaps variants, we conduct FGSM (Goodfellow et al. 2015) and BIM attacks (Kurakin et al. 2017) on the CIFAR-10 dataset. For the iterative attack, BIM is performed for 10 steps. The results are shown in Fig. 4. We can observe a clear distinction between our model variants. For MSPCaps-L, it exhibits exceptional stability, maintaining a high level of accuracy even as the attack strength increases. In contrast, due to parameter constraints, MSPCaps-T’s performance degrades at $\epsilon>0.01$, falling below the original CapsNet’s accuracy.
@@ Conclusion
In this paper, we propose a novel multi-scale patchify capsule network with cross-agreement routing, named MSPCaps. We first introduce MSRB and PatchifyCaps to obtain primary capsules from multi-scale features. Then, to fuse cross-scale capsules effectively for voting, we design a CAR block to adaptively select the most coherent capsules for voting. Experimental results demonstrate that MSPCaps outperforms state-of-the-art methods in terms of classification accuracy and model robustness within capsule networks.
Acknowledgments
This work was supported in part by the Guangdong S&T Programme under Grant No. 2024B0101030002, the Basic Research Project of Hetao Shenzhen-HK S&T Cooperation Zone under Grant No. HZQB-KCZYZ-2021067, the National Natural Science Foundation of China under Grant No. 62501514, the Shenzhen Outstanding Talents Training Fund under Grant No. 202002, the Guangdong Research Projects under Grant No. 2017ZT07X152 and Grant No. 2019CX01X104, the Guangdong Provincial Key Laboratory of Future Networks of Intelligence under Grant No. 2022B1212010001, and the Shenzhen Key Laboratory of Big Data and Artificial Intelligence under Grant No. ZDSYS201707251409055.